Could you show me how to work out this word problem: A shopping center has a rectangular area of 40,000 square yards enclosed on 3 sides for a parking lot. The length is 200 yards more than twice the width. What are the dimensions of the lot? Thanks, Tyler
Answer by checkley77(12844)

L=2W+200
W(2W+200)=40,000
2W^2+200W-40,000=0
2(W^2+100W-20,000)=0
2(W+200)(W-100)=0
W-100=0
W=100 ANSWER FOR THE WIDTH.
L=2*100+200
L=200+200
L=400 ANSWER FOR THE LENGTH.
100*400=40,000
40,000=40,000